Particles moves along the x-axis according to the equation x= 3t^2 - 2t + 3, where x is in meters and t is in seconds at t= 3s. Find a.) the position of the particle b.) it's velocity c.) it's acceleration
1
Expert's answer
2020-09-14T10:20:47-0400

a) The position of the particle at "t = 3s":


"x = 3t^2 - 2t + 3 = 3\cdot (3)^2 - 2\cdot 3 + 3 = 24 m"

b) The velocity of the particle:


"v = \dot{x} = 6t - 2 = 6\cdot 3-2 = 16 m/s"

c)The acceleration of the particle


"a = \dot{v} = 6m/s^2"

Answer. a) 24 m, b) 16 m/s, c) 6 m/s^2.